Help with b and c with explanations please
Question 5. There are 36 students in our class. Suppose I assign everyone a grade of A, B, C, D, or F random. (Really makes you want to study, doesnt it? I promise not to actually assign grades randomly. (a) How many possible assignments of grades are there? 35.34.33.32 30 35.24.3 22 5+.3.2. 86 LA (b) What is the probability that exactly 15 students get an A? (c) What is the probability that YOU get an A? 2
Answer #1

(a) Each student can be assigned any of the 5 grades. So there are 5^36 ways of assigning grades.

(b) 15 students can be selected out of 36 students to assign a grade in C(36, 15) ways = 5567902560 ways, and a grade can be selected to be assigned in 5 ways

Probability = 5567902560 / (5 * 5^36) = 5567902560 / 5^37

(c) There are 5 grades that I can get. So, probability that I get an A grade = 1/5